# Scalable learning of macroscopic stochastic dynamics

### v1

@@ 2.1 Closure modeling of macroscopic dynamics
Eq.˜13 will hold exactly when the macroscopic observable can be expressed as the mean of a function that depends only on individual lattice sites, such as the magnetization. For more complex macroscopic observables, the equality is only approximate, but the approximation improves and becomes accurate in the limit of a large system size.
@@
If $\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}$ is independent of $\mathbf{z}_{t+\delta t,\mathcal{J}}-\mathbf{z}_{t}$ when $\mathcal{I}\neq\mathcal{J}$, and if both Eq.˜12 and Eq. 13 hold exactly:
$\mathbf{x}_{t+\delta t,\mathcal{I}}$ $\overset{d}{=}\tilde{\mathbf{x}}_{t+\delta t,\mathcal{I}},\quad\forall I\in\mathcal{I},$ (14)
$\boldsymbol{\varphi}^{\ast}(\mathbf{x})$ $=\tfrac{1}{K}{\textstyle\sum_{\mathcal{I}}}\boldsymbol{\varphi}^{\ast}(\mathbf{x}_{\mathcal{I}}).$
Then there exists a unique minimizer $(\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast})$ of $\mathcal{L}\allowbreak[\boldsymbol{\mu},\boldsymbol{\Sigma}]$, and a unique minimizer $(\boldsymbol{\mu}^{\dagger},\boldsymbol{\Sigma}^{\dagger})$ of $\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$, such that:
$\boldsymbol{\mu}^{\ast}=\boldsymbol{\mu}^{\dagger},\quad\boldsymbol{\Sigma}^{\ast}=\boldsymbol{\Sigma}^{\dagger}.$ (15)
The proof is based on a direct computation of the first and second variations of the loss functions. We provide the full proof in Appendix˜B. Theorem˜1 theoretically justifies our framework: Under appropriate conditions, the macroscopic dynamics learned from data generated via partial evolution are as accurate as those learned from full, computationally expensive microscopic simulations.
2.2 Hierarchical upsampling scheme
@@
4:Initialize $\mathcal{D}^{(0)}\leftarrow\mathcal{D}_{s}$
@@
6:  $\mathcal{D}^{(i)}\leftarrow\textsc{LocalRelax}\big(\textsc{Upsample}(\mathcal{D}^{(i-1)})\big)$
@@
8:return $\mathcal{D}$ $\triangleright$ data distribution of the large system
In Section˜2.1, we assume access to the data distribution $\mathcal{D}$, which contain multiple snapshots of large systems. In practice, however, direct access to $\mathcal{D}$ is typically unavailable, as only microscopic simulations of small systems can be performed. To address this, We introduce a hierarchical upsampling scheme for generating the large-system data distribution $\mathcal{D}$ from the small-system trajectory distribution $\mathcal{D}_{s}$. The hierarchical upsampling scheme is illustrated in Algorithm˜1. It consists of multiple iterations, each involving two steps: Upsample and LocalRelax. In the Upsample step, the configurations in $\mathcal{D}^{(i)}$ are expanded into configurations of size $m$ times larger, where $m\geq 2$ is an integer. Next, we apply a LocalRelax step to remove the unphysical artifacts that are introduced in the Upsample step. More specifically, each generated large-system configuration is divided into overlapping patches of size $n_{s}$, and short-time relaxation or local dynamics evolution is applied within each patch.
For example, consider the two-dimensional Ising Model to be introduced in Section˜3.3, where the microscopic dynamics is chosen to be the continuous-time Glauber dynamics. We assume direct simulations are feasible only for systems up to size $n_{s}=8^{2}$. Starting from the small-system trajectory distribution $\mathcal{D}_{s}$, we apply the hierarchical upsampling scheme for $N_{\text{iter}}=3$ iterations to obtain the large-system data distribution $\mathcal{D}$ of size $n=64^{2}$. In the first iteration, the upsample step replicates each spin into a $m=2^{2}$ block, yielding a $16^{2}$ dimensional system. Next, we apply the LocalRelax step by dividing each $16^{2}$ dimensional configuration into $16$ patches of size $n_{s}=8^{2}$ with a stride of $4$. Within each patch, we run the continuous-time Glauber dynamics for a short time to remove the unphysical artifacts introduced in the upsample step. In the second iteration, starting from the $16^{2}$ dimensional data distribution $\mathcal{D}^{(1)}$, we repeat the upsample and LocalRelax step to obtain $\mathcal{D}^{(2)}$ for the $32^{2}$ dimensional system. Subsequently, in the third iteration, we obtain the target large-system distribution $\mathcal{D}$ of size $64^{2}$ from $\mathcal{D}^{(2)}$. We provide a graphical illustration of one iteration for the Ising model in Fig.˜1 (a). The concrete form of Upsample and LocalRelax may vary for different microscopic systems, and further details are given in Section˜3.
@@ 3 RESULTS
3.1 Stochastic Predator-Prey system
@@
We first consider a one-dimensional SPDE system, mainly to validate the correctness of our method and investigate the impact of the coefficient $K$ in the designed loss $\mathcal{L}_{p}$. The stochastic Predator-Prey system is given by Eq.˜16, where $u,v$ denote the dimensionless populations of the prey and predator, and $\xi_{u},\xi_{v}$ are independent space-time white noise terms. In our experiment, we set the parameters $a=3,b=0.4,D=0$ and $\sigma_{u}=\sigma_{v}=0.02$.
$\frac{\partial u}{\partial t}=u(1-u-v)+D\frac{\partial^{2}u}{\partial x^{2}}+\sigma_{u}\,\xi_{u}(t,x),$ (16)
@@
For training data generation, we first generate the small-system distribution $\mathcal{D}_{s}$ by discretizing the spatial domain $\Omega$ into $100$ uniform grids. We then solve Eq.˜16 from $t=0$ to $T=30$ with time step $\delta t=0.01$. The small system thus contains $n_{s}=100$ lattice sites. The goal is to learn the macroscopic observables of a large system discretized on $n=200$ uniform grid points. To construct the large-system distribution $\mathcal{D}$, we perform an Upsample step by linearly interpolating the solution from the coarse grid with $100$ points onto a finer grid with $200$ points. Since this interpolation does not introduce unphysical artifacts, the LocalRelax step is not required here. The $n=200$ uniform grids are partitioned into $K=5$ patches, each containing $40$ grids. For each $\mathbf{x}_{t}\sim\mathcal{D}$, we first uniformly sample a patch with $p=1/5$, and subsequently evolve the system locally within the patch for one time step $\delta t=0.01$ to obtain the updated state $\mathbf{x}_{t+\delta t,\mathcal{I}}$.
@@
The macroscopic observable $\mathbf{z}^{\ast}$ of interest is chosen to be the mean of the populations of the prey and predator over the spatial grid points, which is two-dimensional. We identify another 2 closure variables using an autoencoder, hence the dimension of $\mathbf{z}$ is 4. The only difference between $\mathcal{L}$ and $\mathcal{L}_{p}$ lies in the coefficient multiplying the diffusion term. We treat the coefficient as a hyperparameter $\lambda$ and train the SDE with different values of $\lambda$. Theoretical analysis in Theorem˜1 shows that the optimal value of $\lambda$ is $K$ under appropriate conditions. We explore various values of $\lambda$ to empirically investigate the influence of $\lambda$.
@@
Fig.˜2 shows the result on the $9$ test datasets with different combinations of parameters $(c_{1},c_{2})\in\{0.05,\allowbreak 0.10,0.15\}\times\{0.45,0.50,0.55\}$. Each test dataset consists of $50$ trajectories with the same initial condition. Once the SDE model in Eq.˜4 is trained, we simulate it for a long time with the Euler-Maruyama method starting from the same initial condition as the test dataset. We employ Maximum Mean Discrepancy (MMD) [31, 32] to quantify the discrepancy between the predicted and ground-truth trajectories, which is widely used for comparing probability distributions. We use a mixture of radial basis function (RBF) kernels with varying bandwidths to improve the robustness of MMD. For each time point, we calculate the MMD between the marginal distributions of the predicted trajectories and the ground truth trajectories, and report the results averaged over all $9$ test datasets.
@@
From Fig.˜2 we can observe that when $\lambda=K=5$, the predicted trajectories achieve the minimal discrepancy from the ground truth trajectories, which correspond well with our theoretical analysis. In fact, for the stochastic Predator-Prey system, the assumptions in Theorem˜1 are exactly satisfied. Therefore, it is expected that the optimal hyperparameter $\lambda$ is equal to $K$. In subsequent experiments with more complex microscopic dynamics, the assumptions in Theorem˜1 may only hold approximately. Thus, the optimal value of $\lambda$ may deviate slightly from $K$. We will treat $\lambda$ as a tunable hyperparameter and perform a hyperparameter search initialized from $K$ in the following experiments.
3.2 Curie-Weiss model
Having validated our method on a toy SPDE system, we now turn to more complex spin systems. We first consider the Curie-Weiss model, which is a mean-field approximation of the Ising model. The Curie-Weiss model is widely used in statistical mechanics to describe phase transitions due to its simplicity [33, 34]. In this experiment, we conduct an ablation study on the computational power $n_{s}$ of the microscopic simulator $\mathcal{S}_{n_{s}}$, and compare the performance of our method with the baseline across different values of $n_{s}$.
The Hamiltonian of the Curie-Weiss Model with $n$ spins is given by:
$H_{n,h}(\sigma)=-\frac{J}{2n}\sum_{i,j=1}^{n}\sigma_{i}\sigma_{j}-h\sum_{i=1}^{n}\sigma_{i},$ (19)
where $J>0$ is the interaction strength, $h$ is the external magnetic field and $\sigma_{i}\in\{-1,1\}$ denotes the spin state. Throughout the remainder of this paper, we adopt dimensionless units by setting $J=1$ and the Boltzmann constant $k_{B}=1$. Eq.˜19 can also be expressed as a function of the magnetization $M=\sum_{i}\sigma_{i}/n$:
$H_{n,h}(\sigma)=-n\left[\frac{J}{2}\left(\frac{\sum_{i}\sigma_{i}}{n}\right)^{2}+h\frac{\sum_{i}\sigma_{i}}{n}\right].$ (20)
Hence, the microscopic dynamics can be fully characterized by the magnetization. We choose the magnetization as the macroscopic observable of interest since the magnetization is closed by itself, and no additional closure variables are needed.
Strictly speaking, the Curie-Weiss model does not fit directly into our framework, because it is a mean-field model without any explicit spatial structure. To handle this, we imagine the spins are arranged on a square lattice. Moreover, the interaction between spins is nonlocal since each spin interacts with all others through the global coupling. To overcome this, here we take a sufficiently large local patch of size $n_{s}$, such that the magnetization computed within the local patch can accurately represent the global magnetization.
@@ 3.2 Curie-Weiss model
Figure 3: Test error as a function of $n_{s}$. The test error is the mean relative error of the mean macroscopic observables between ground-truth and predicted trajectories. (a) Results on the Curie-Weiss model. (b) Results on the Ising model.
Figure 4: Results on the Curie-Weiss model with $n_{s}=16^{2}$. Mean and standard deviation of the magnetization are estimated from 2020 trajectories per method.
@@
Figure 5: Results on the Ising model with $n_{s}=16^{2}$. Mean and standard deviation are estimated from 2020 trajectories per method. (a) Magnetization statistics. (b) Domain wall density statistics.
In this experiment, the microscopic state is $\mathbf{x}=\{\sigma_{1},\cdots,\sigma_{n}\}$. We fix the large-system size $n=64^{2}$, while varying the computational power of the microscopic simulator $\mathcal{S}_{n_{s}}$ by considering $n_{s}\in\{8^{2},16^{2},32^{2},64^{2}\}$. When $n_{s}=64^{2}$, our method reduces to the traditional methods for identifying the macroscopic dynamics. The macroscopic observable $\mathbf{z}^{\ast}$ is chosen to be the magnetization $\sum_{i}\sigma_{i}/n$. No closure variable is needed, and the latent dimension is $1$. We consider the microscopic dynamics to be the continuous-time Glauber dynamics [35], as described in Algorithm˜2. At each step, the energy differences $\{\Delta H_{i}\}_{i=1}^{n}$ and flip rates $\{r_{i}\}_{i=1}^{n}$ need to be calculated for all spins.
We generate the small-system distribution $\mathcal{D}_{s}$ by simulating the continuous-time Glauber dynamics with $h=0.1$ at temperature $T=1.1$, which is slightly higher than the critical temperature $T_{c}=1$. To generate the large-system data distribution $\mathcal{D}$, we employ the hierarchical algorithm with $\log_{2}(n/n_{s})$ iterations. For each iteration, we first perform an Upsample step to replicate every spin into $2^{2}$ block. Next, we perform a LocalRelax step by evolving a short-time continuous-time Glauber dynamics locally to remove the unphysical artifacts introduced in the upsample step.
Fig.˜3 (a) and Fig.˜4 compare the performance of our method with the baseline. From Fig.˜3 (a), we observe that our method consistently outperforms the baseline when $n_{s}<n$, which demonstrates the effectiveness of our method. Fig.˜4 further compares the magnetization statistics derived from ground truth trajectories with those produced by the baseline model and our method. The statistics of the trajectories predicted by our method align closely with the ground truth, whereas the baseline exhibits much larger variations.
The parameter $n_{s}$ can influence our method in two ways: (i) The size of the small-system distribution $\mathcal{D}_{s}$ is $n_{s}$. For smaller $n_{s}$, more iterations of upsampling are required to obtain $\mathcal{D}$, which may degrade the data quality of $\mathcal{D}$. (ii) When generating $\mathbf{x}_{t+\delta t,\mathcal{I}}$ from $\mathbf{x}\sim\mathcal{D}$, we perform partial evolution on a small patch of size $n_{s}$. A smaller $n_{s}$ leads to higher stochasticity of the generated data. As a result, a smaller $n_{s}$ generally leads to worse performance. The test error of our method remains relatively low when $n_{s}\geq 16^{2}$, which demonstrates the robustness of our method w.r.t. $n_{s}$.
3.3 Ising model
@@
Figure 6: Finite size scaling results of the Ising model at $T=2.27$, which is very close to $T_{c}\approx 2.269$. (a) Scaling of the equilibrium magnetization, with a fitted value $(\beta/\nu)^{\ast}=0.120$. (b) Scaling of the magnetic susceptibility, with a fitted value $(\gamma/\nu)^{\ast}=1.792$. (c) Raw Binder cumulant data for different temperatures and linear size $L$. Linear fitting is used for $L=16$, and cubic polynomial fitting for $L=32,48,64$. The slope at $T_{c}$ is extracted from the fitted curve. (d) Log-log plot of $|dU/dT|_{T=T_{c}}$ versus $L$, with a fitted value $\nu^{\ast}=1/1.05\approx 0.95$.
@@
We next validate our method on the Ising model, a more realistic setting than the Curie–Weiss model because it incorporates local interactions and spatial structure. The Ising model plays an important role in statistical physics for investigating order-disorder phase transitions and critical phenomena [36, 37]. We will validate the capability of our method to accurately identify the critical behavior of macroscopic dynamics.
@@
$H_{n,h}(\sigma)=-\frac{J}{2}\sum_{<i,j>}\sigma_{i}\sigma_{j}-h\sum_{i=1}^{n}\sigma_{i},$ (21)
where $\langle i,j\rangle$ denotes the first nearest neighbor. As in the Curie–Weiss model, we adopt the continuous-time Glauber dynamics as the microscopic dynamics. The critical temperature of the two-dimensional Ising model is $T_{c}\approx 2.269$. The Ising model exhibits an ordered ferromagnetic phase when $T<T_{c}$, and a disordered paramagnetic phase when $T>T_{c}$.
@@
The domain wall density quantifies the degree of disorder by measuring the proportion of misaligned nearest-neighbor spin pairs. We identify another $2$ closure variables by the autoencoder, resulting in a latent state $\mathbf{z}$ of dimension $4$.
The experiments are divided into two parts. In the first part, we repeat the procedure of Section˜3.2 for the Ising model with $h=0.1,T=2.5>T_{c}$. We employ the same Upsample and LocalRelax step as used in the Curie-Weiss model for data generation. The results are shown in Fig.˜3(b) and Fig.˜5. It can be observed that our method achieves lower test error than the baseline. In particular, when $n_{s}=8^{2}$, the test error of the baseline increases significantly, while the test error remains relatively low. Fig.˜5 demonstrates that our method can accurately reproduce the statistics of trajectory distributions, while the predicted trajectories of the baseline exhibit much higher variance.
  (a) Microscopic Configuration (b) Order Parameters (c) Macroscopic Dynamics
@@
In Fig.˜6 (a) and (b), we plot the equilibrium magnetization and susceptibility as a function of $L$ at $T=2.27$, which is very close to $T_{c}$. Next, we fit log-log curves to the data by Eq.˜25, yielding estimated critical exponent $(\beta/\nu)^{\ast}=0.120,(\gamma/\nu)^{\ast}=1.792$. The estimated critical exponents are very close to the theoretical value $\beta/\nu=0.125,\gamma/\nu=1.75$. These results show that our method can accurately recover the ratios $\beta/\nu$ and $\gamma/\nu$, demonstrating its ability to capture accurate macroscopic dynamics across system sizes. This is especially significant because we only use microscopic simulations of small systems with size $n_{s}=16^{2}$, yet our method still succeeds in reproducing macroscopic dynamics for large systems with different sizes.

### v2

@@ II.1 Closure modeling of macroscopic dynamics
Similarly, the macroscopic increment $\Delta\mathbf{z}^{\ast}_{t}=\boldsymbol{\varphi}^{\ast}(\mathbf{x}_{t+\delta t})-\boldsymbol{\varphi}^{\ast}(\mathbf{x}_{t})$ can be approximated by the average of the local increments,
$\Delta\mathbf{z}^{\ast}_{t}\approx\frac{1}{K}\sum_{\mathcal{I}}\Delta\mathbf{z}^{\ast}_{t,\mathcal{I}}.$ (15)
Eqs. 14 and 15 hold exactly when the macroscopic observable can be written as the mean of a function depending only on individual lattice sites, such as the magnetization. For more complex macroscopic observables, the equality is only approximate, but the approximation improves and becomes accurate in the limit of a large system size.
@@
Assume that for any two disjoint patches $\mathcal{I}\neq\mathcal{J}$, the short-time increments $\Delta\mathbf{z}_{t,\mathcal{I}}$ and $\Delta\mathbf{z}_{t,\mathcal{J}}$ are conditionally independent given $\mathbf{x}_{t}$. Assume further that the encoder $\boldsymbol{\varphi}$ is uniformly bounded, i.e., $\|\boldsymbol{\varphi}\|_{\infty}\leq M$ for some $M>0$.
Suppose that Eqs. 13 and 15 hold up to second-order accuracy in $\delta t$, in the sense that the total variation distance between $q$ and $\hat{q}$ satisfies
$\delta_{\mathrm{TV}}\!\left(q(\cdot\mid\mathbf{x}_{t}),\hat{q}(\cdot\mid\mathbf{x}_{t})\right)\leq C_{1}\delta t^{2},$ (16)
and that the macroscopic increment approximation error satisfies
$\bigl\|\mathbb{E}_{\mathbf{x}_{t+\delta t}\mid\mathbf{x}_{t}}[\Delta\mathbf{z}_{t}^{\ast}-\tfrac{1}{K}\sum_{\mathcal{I}}\Delta\mathbf{z}^{\ast}_{t,\mathcal{I}}]\bigr\|\leq C_{2}\delta t^{2},$ (17)
where $C_{1},C_{2}>0$ are constants and $\mathbf{x}_{t+\delta t}\mid\mathbf{x}_{t}$ denotes the ground-truth conditional distribution obtained by evolving the full system from $\mathbf{x}_{t}$ for time $\delta t$.
Under these assumptions, the functional $\mathcal{L}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ admits a unique minimizer $(\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast})$, and the functional $\mathcal{L}_{p}[\boldsymbol{\mu},\boldsymbol{\Sigma}]$ admits a unique minimizer $(\boldsymbol{\mu}^{\dagger},\boldsymbol{\Sigma}^{\dagger})$, such that
$\|\boldsymbol{\mu}^{\ast}-\boldsymbol{\mu}^{\dagger}\|_{\infty}$ $\leq K_{1}\delta t,$ (18)
$\|\boldsymbol{\Sigma}^{\ast}-\boldsymbol{\Sigma}^{\dagger}\|_{\infty}$ $\leq K_{2}\delta t,$
where $K_{1},K_{2}>0$ are constants depending only on $C_{1}$, $C_{2}$, and $M$. In particular, if $q=\hat{q}$ holds exactly, we have
$\boldsymbol{\mu}^{\ast}=\boldsymbol{\mu}^{\dagger},\qquad\boldsymbol{\Sigma}^{\ast}=\boldsymbol{\Sigma}^{\dagger}.$ (19)
We emphasize that the independence assumption in Theorem 1 is about short-time increments on disjoint patches conditioned on $\mathbf{x}_{t}$, not about independence of the patch states themselves. The proof is based on a direct computation of the first and second variations of the loss functions. We provide the full proof in Appendix C. Theorem 1 theoretically justifies our framework: Under appropriate conditions, the macroscopic dynamics learned from data generated via partial evolution are as accurate as those learned from full, computationally expensive microscopic simulations. In particular, for the stochastic Predator–Prey system in Section III.1, the assumptions of Theorem 1 are satisfied, as verified in Appendix E.
II.2 Hierarchical upsampling scheme
@@
4:Initialize $D^{(0)}\leftarrow D_{s}$
@@
6:  $D^{(i)}\leftarrow\textsc{LocalRelax}\big(\textsc{Upsample}(D^{(i-1)})\big)$
@@
8:return $D$ $\triangleright$ dataset of large-system snapshots
In Section II.1, we assume access to the dataset $D$, which contains multiple snapshots of large systems. In practice, however, direct access to $D$ is typically unavailable, as only microscopic simulations of small systems can be performed. To address this, we introduce a hierarchical upsampling scheme for generating the large-system dataset $D$ from the small-system dataset $D_{s}$. The hierarchical upsampling scheme is illustrated in Algorithm 1. It consists of multiple iterations, each involving two steps: Upsample and LocalRelax. In the Upsample step, the configurations in $D^{(i)}$ are expanded into configurations of size $m$ times larger, where $m\geq 2$ is an integer. Next, we apply a LocalRelax step to remove the unphysical artifacts that are introduced in the Upsample step. More specifically, each generated large-system configuration is divided into overlapping patches of size $n_{s}$, and short-time relaxation or local dynamics evolution is applied within each patch.
For example, consider the two-dimensional Ising Model to be introduced in Section III.2, where the microscopic dynamics is chosen to be the continuous-time Glauber dynamics. We assume direct simulations are feasible only for systems up to size $n_{s}=8^{2}$. Starting from the small-system dataset $D_{s}$, we apply the hierarchical upsampling scheme for $N_{\text{iter}}=3$ iterations to obtain the large-system dataset $D$ of size $n=64^{2}$. In the first iteration, the upsample step replicates each spin into a $m=2^{2}$ block, yielding a $16^{2}$ dimensional system. Next, we apply the LocalRelax step by dividing each $16^{2}$ dimensional configuration into $16$ patches of size $n_{s}=8^{2}$ with a stride of $4$. Within each patch, we run the continuous-time Glauber dynamics for a short time to remove the unphysical artifacts introduced in the upsample step. In the second iteration, starting from the $16^{2}$ dimensional dataset $D^{(1)}$, we repeat the upsample and LocalRelax step to obtain $D^{(2)}$ for the $32^{2}$ dimensional system. Subsequently, in the third iteration, we obtain the target large-system dataset $D$ of size $64^{2}$ from $D^{(2)}$. We provide a graphical illustration of one iteration for the Ising model in Fig. 1 (a). The concrete form of Upsample and LocalRelax may vary for different microscopic systems, and further details are given in Section III.
@@ III RESULTS
III.1 Stochastic Predator-Prey system
@@
We first consider a one-dimensional SPDE system, mainly to validate the correctness of our method and investigate the impact of the coefficient $K$ in the designed loss $\mathcal{L}_{p}$. The stochastic Predator-Prey system is given by Eq. 20, where $u,v$ denote the dimensionless populations of the prey and predator, and $\xi_{u},\xi_{v}$ are independent space-time white noise terms. In our experiment, we set the parameters $a=3,b=0.4,c=0$ and $\sigma_{u}=\sigma_{v}=0.02$.
$\frac{\partial u}{\partial t}=u(1-u-v)+c\frac{\partial^{2}u}{\partial x^{2}}+\sigma_{u}\,\xi_{u}(t,x),$ (20)
@@
For training data generation, we first generate the small-system dataset $D_{s}$ by discretizing the spatial domain $\Omega$ into $100$ uniform grids. We then solve Eq. 20 from $t=0$ to $T=30$ with time step $\delta t=0.01$. The small system thus contains $n_{s}=100$ lattice sites. The goal is to learn the macroscopic observables of a large system discretized on $n=200$ uniform grid points. To construct the large-system dataset $D$, we perform an Upsample step by linearly interpolating the solution from the coarse grid with $100$ points onto a finer grid with $200$ points. Compared to snapshots sampled from large-system trajectories, snapshots generated by linear interpolation are smoother. However, we expect the resulting differences to be small. Therefore, for simplicity, we do not apply the LocalRelax step.
The $n=200$ uniform grids are partitioned into $K=5$ patches, each containing $40$ grids. For each $\mathbf{x}_{t}\sim D$, we first uniformly sample a patch with $p=1/5$, and subsequently evolve the system locally within the patch for one time step $\delta t=0.01$ to obtain the updated state $\mathbf{x}_{t+\delta t,\mathcal{I}}$.
@@
The macroscopic observable $\mathbf{z}^{\ast}$ of interest is chosen to be the mean of the populations of the prey and predator over the spatial grid points, which is two-dimensional. We derive another 2 closure variables using an autoencoder, hence the dimension of $\mathbf{z}$ is 4. The only difference between $\mathcal{L}$ and $\mathcal{L}_{p}$ lies in the coefficient multiplying the diffusion term. We treat the coefficient as a hyperparameter $\lambda$ and train the SDE with different values of $\lambda$. Theoretical analysis in Theorem 1 shows that the optimal value of $\lambda$ is $K$ under appropriate conditions. We explore various values of $\lambda$ to empirically investigate the influence of $\lambda$.
@@
Fig. 2 shows the result on the $9$ test datasets with different combinations of parameters $(c_{1},c_{2})\in\{0.05,\allowbreak 0.10,0.15\}\times\{0.45,0.50,0.55\}$. Each test dataset consists of $50$ trajectories with the same initial condition. Once the SDE model in Eq. 4 is trained, we simulate it for a long time with the Euler-Maruyama method starting from the same initial condition as the test dataset. We employ Maximum Mean Discrepancy (MMD) [30, 35] to quantify the discrepancy between the predicted and ground-truth trajectories, which is widely used for comparing probability distributions. We use a mixture of radial basis function (RBF) kernels with varying bandwidths to improve the robustness of MMD. For each time point, we calculate the MMD between the marginal distributions of the predicted trajectories and the ground truth trajectories, and report the results averaged over all $9$ test datasets.
@@
From Fig. 2 we can observe that when $\lambda=K=5$, the predicted trajectories achieve the minimal discrepancy from the ground truth trajectories, which correspond well with our theoretical analysis. In fact, for the stochastic Predator-Prey system, the assumptions in Theorem 1 are exactly satisfied. Therefore, it is expected that the optimal hyperparameter $\lambda$ is equal to $K$. In subsequent experiments with more complex microscopic dynamics, the assumptions in Theorem 1 may only hold approximately. Thus, the optimal value of $\lambda$ may deviate slightly from $K$. We will treat $\lambda$ as a tunable hyperparameter and perform a hyperparameter search initialized from $K$ in the following experiments.
III.2 Ising model
@@
Figure 3: Results on the Ising model, where the test error is plotted as a function of $n_{s}$. The test error is the mean relative error of the mean macroscopic observables between ground-truth and predicted trajectories.
@@
Figure 4: Results on the Ising model with $n_{s}=16^{2}$. Mean and standard deviation are estimated from $20$ trajectories per method. (a) Magnetization statistics. (b) Domain wall density statistics.
@@
Figure 5: Finite size scaling results of the Ising model at $T=2.27$, which is very close to $T_{c}\approx 2.269$. (a) Scaling of the equilibrium magnetization, with a fitted value $(\beta/\nu)^{\ast}=0.129$. (b) Scaling of the magnetic susceptibility, with a fitted value $(\gamma/\nu)^{\ast}=1.670$. (c) Raw Binder cumulant data for different temperatures and linear size $L$. Linear fitting is used for $L=16$, and cubic polynomial fitting for $L=32,48,64$. The slope at $T_{c}$ is extracted from the fitted curve. (d) Log-log plot of $|dU/dT|_{T=T_{c}}$ versus $L$, with a fitted value $\nu^{\ast}=1/1.035\approx 0.97$.
@@
Having validated our method on a toy SPDE system, we now turn to more complex spin systems. We first consider the Ising model, which plays an important role in statistical physics for investigating order-disorder phase transitions and critical phenomena [7, 14]. The experiments are divided into two parts. In the first part, we conduct an ablation study on the computational power $n_{s}$ of the microscopic simulator $\mathcal{S}_{n_{s}}$, and compare the performance of our method with the baseline across different values of $n_{s}$. In the second part, we further evaluate the ability of our method to accurately capture the critical behavior of macroscopic dynamics.
@@
$H_{n,h}(\sigma)=-\frac{J}{2}\sum_{\langle i,j\rangle}\sigma_{i}\sigma_{j}-h\sum_{i=1}^{n}\sigma_{i},$ (23)
where $J>0$ denotes the interaction strength, $h$ denotes the external magnetic field, $\langle i,j\rangle$ represents nearest-neighbor pairs, and $\sigma_{i}\in\{-1,1\}$ denotes the spin at site $i$. The microscopic state is denoted by $\mathbf{x}=\{\sigma_{1},\cdots,\sigma_{n}\}$. Throughout this paper, we employ dimensionless units by setting $J=1$ and the Boltzmann constant $k_{B}=1$. We describe the microscopic evolution of the system using continuous-time Glauber dynamics, as detailed in Appendix B. The critical temperature of the two-dimensional Ising model is $T_{c}\approx 2.269$. The Ising model exhibits an ordered ferromagnetic phase when $T<T_{c}$, and a disordered paramagnetic phase when $T>T_{c}$.
@@
The domain wall density quantifies the degree of disorder by measuring the fraction of misaligned nearest-neighbor spin pairs. We derive two additional closure variables using an autoencoder, resulting in a latent state $\mathbf{z}$ of dimension $4$.
In the first part, we fix the large-system size to $n=64^{2}$, while varying the computational power of the microscopic simulator $\mathcal{S}_{n_{s}}$ by considering $n_{s}\in\{8^{2},16^{2},32^{2},64^{2}\}$. When $n_{s}=64^{2}$, our method reduces to traditional methods for deriving the macroscopic dynamics. We set the external field to $h=0.1$ and the temperature to $T=2.5>T_{c}$. To generate the large-system dataset $D$, we employ the hierarchical algorithm with $\log_{2}(n/n_{s})$ iterations. For each iteration, we first perform an Upsample step to replicate every spin into $2^{2}$ block. Next, we perform a LocalRelax step by evolving a short-time continuous-time Glauber dynamics locally to remove the unphysical artifacts introduced in the Upsample step.
Fig. 3 and Fig. 4 compare the performance of our method with the baseline. From Fig. 3, we observe that our method consistently outperforms the baseline when $n_{s}<n$, which demonstrates the effectiveness of our method. Fig. 4 further compares the magnetization statistics and domain wall density statistics derived from ground truth trajectories with those produced by the baseline model and our method. The statistics of the trajectories predicted by our method align closely with the ground truth, whereas the baseline exhibits much larger variations.
The parameter $n_{s}$ can influence our method in two ways: (i) The size of the small-system dataset $D_{s}$ is $n_{s}$. For smaller $n_{s}$, more iterations of upsampling are required to obtain $D$, which may degrade the data quality of $D$. (ii) When generating $\mathbf{x}_{t+\delta t,\mathcal{I}}$ from $\mathbf{x}\sim D$, we perform partial evolution on a small patch of size $n_{s}$. A smaller $n_{s}$ leads to higher stochasticity of the generated data. As a result, a smaller $n_{s}$ generally leads to worse performance. However, the test error of our method remains relatively low when $n_{s}\geq 16^{2}$, which demonstrates the robustness of our method w.r.t. $n_{s}$.
  (a) Microscopic Configuration (b) Order Parameters vs. Temperature (c) Macroscopic Dynamics
@@
In Fig. 5 (a) and (b), we plot the equilibrium magnetization and susceptibility as a function of $L$ at $T=2.27$, which is very close to $T_{c}$. Next, we fit log-log curves to the data by Eq. 27, yielding estimated critical exponent $(\beta/\nu)^{\ast}=0.129,(\gamma/\nu)^{\ast}=1.670$. The estimated critical exponents are very close to the theoretical value $\beta/\nu=0.125,\gamma/\nu=1.75$. These results show that our method can accurately recover the ratios $\beta/\nu$ and $\gamma/\nu$, demonstrating its ability to capture accurate macroscopic dynamics across system sizes. This is especially significant because we only use microscopic simulations of small systems with size $n_{s}=16^{2}$, yet our method still succeeds in reproducing macroscopic dynamics for large systems with different sizes.